OpenStudy (narissa):

The length of a rectangle is 3 more than 4 times its width. The perimeter of the rectangle is 86 inches. What is the length of the rectangle?

OpenStudy (narissa):

@AlexandervonHumboldt2

OpenStudy (narissa):

@triciaal

OpenStudy (narissa):

@wwhitlock

OpenStudy (narissa):

@CShrix

OpenStudy (wwhitlock):

4w + 3 = l w + w + l + l = 86

OpenStudy (narissa):

so i dk do i add all of those or what?

OpenStudy (narissa):

is it 27?

OpenStudy (wwhitlock):

substitute 4w + 3 for each time you have an l in the second line.

OpenStudy (narissa):

23?

OpenStudy (wwhitlock):

2w + 2(4w + 3) = 86

OpenStudy (narissa):

35?

OpenStudy (narissa):

or 31?

OpenStudy (narissa):

did I pick one of the correct answers or no?

OpenStudy (wwhitlock):

lots of guessing. What's the width? and what's the length? 23 is one of them

OpenStudy (narissa):

23 is the length lol

OpenStudy (narissa):

thanks i was in a hurry i put 23 as the length but oh well

OpenStudy (narissa):

got it wrong

OpenStudy (wwhitlock):

oops so did I. Trying again.

OpenStudy (narissa):

it okay i cant do it again

OpenStudy (wwhitlock):

sorry about that. I skipped a step. I'm better at history than math.

OpenStudy (wwhitlock):

you should take the medal, though.